# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ IV.2 Mode amplitude validation
We now turn to examine the faithfulness of the amplitude module of our waveform generation framework. In general, any systematic errors in the mode amplitudes $A_{\ell mn}$ should be significantly less impactful than those present in the forcing functions. Contributions from the latter grow secularly over inspiral, whereas amplitude errors remain of fixed order. For systematic biases in EMRI parameter estimation to be sufficiently small (assuming no orbital phase errors), the relative error in each mode amplitude must be smaller (in an order-of-magnitude sense) than the inverse of the SNR of the mode lindblom2008model . It is therefore important to verify that the scale of these errors is small. In this subsection, we will compare the outputs of our amplitude module with an independent dataset to quantify the impact of errors in our interpolated amplitudes. We will also explore where our choice to perform linear interpolation with respect to $a$ (as described in Section III.2.2) limits the accuracy our mode amplitude values when compared to spline interpolation along this dimension.
As in Section IV.1, we will validate our model by comparing interpolated mode amplitudes ($\mathcal{A}_{\ell mn}^{\mathrm{FEW}}$) against those produced with the GREMLIN code ($\mathcal{A}_{\ell mn}^{\mathrm{SAH}}$). As our amplitude module produces $6993$ modes at each point in parameter space, examining any particular mode in isolation will not be particularly informative with regards to the accuracy of our amplitude model as a whole (especially as the relative importance of each mode is also a function of the parameter space). Instead, in a similar vein to Ref. Isoyama:2021jjd , we will define a suitable figure of merit that encapsulates all mode amplitudes. It is similar in form to that of the waveform mismatch (which we will apply in Section IV.3; see Appendix D for details), and is written as
@@
where $\star$ denotes complex conjugation and ||$\cdot$|| ⋅ | is the amplitude vector norm (over the indices $\ell$, $m$ and $n$). This quantity corresponds to a mismatch between two snapshot waveforms (i.e., constant in amplitude and frequency) with amplitudes $\mathcal{A}^{\mathrm{FEW}}_{\ell mn}$ and $\mathcal{A}^{\mathrm{SAH}}_{\ell mn}$, assuming that overlap terms between modes is small. While this is not generally true, the inclusion of these terms is not required for us to explore how the accuracy of our modes varies over parameter space, and their exclusion enables us to avoid performing waveform computations at this stage. This also ignores the spin-weighted spherical harmonics, as we wish to investigate differences independent of the viewing angle. The $\mathcal{A}^{\mathrm{SAH}}_{\ell mn}$ data we compare against only extends to $p\sim 10$; see Section E.2 for a comparison between our amplitude interpolation and PN results in the weak-field limit. In Fig. 9, we show how $\mathcal{M}_{\mathrm{amp}}$ varies over $(p,e)$ for $a=0.8952$ (which lies between two nodes in our amplitude data grids, maximising interpolation errors). For $p\gtrsim p_{\mathrm{sep}}+1$, our interpolated amplitudes agree well with the GREMLIN data for all eccentricities, with estimated mismatches $\leq 10^{-5}$. Below this value, $\mathcal{M}_{\mathrm{amp}}$ grows rapidly with respect to $p$, reaching $\sim 10^{-3}$ at higher eccentricities for $p\gtrsim p_{\mathrm{sep}}+10^{-2}$. This suggests that our amplitude interpolation errors are significant near the separatrix. In Section IV.3, we investigate the impact of amplitude interpolation on the waveform level and obtain results that agree quantitatively with those presented in Fig. 9.
@@
Figure 9: Mode amplitude (approximate) mismatch $\mathcal{M}_{\mathrm{amp}}$ (Eq. 14) between interpolated amplitudes from our model ($\mathcal{A}_{\ell mn}^{\mathrm{FEW}}$) and those produced with the GREMLIN code ($\mathcal{A}_{\ell mn}^{\mathrm{SAH}}$). Mismatch due to amplitude interpolation error increases significantly as $p$ decreases below $\sim p_{\mathrm{sep}}+1$, and is slightly worse for higher eccentricities. This behaviour is due to our linear interpolation of amplitudes with respect to $a$ (explored in Fig. 10).
@@
Figure 10: Relative difference between bicubic+linear (BIC) and tricubic interpolation (TRI) of the $A_{220}$ mode amplitude with respect to: $(p,e)$ with $a=0.998$ (top panel); $(p,a)$ with grid coordinate $w=0.5$ (bottom panel). The eccentricity corresponding to this value of $w$ for $a=0.998$ is shown in the top panel as a cyan dashed line; for lower spins, $e$ at $p=p_{\mathrm{sep}}+10^{-3}$ will be higher than this line (see Appendix B). The outputs of the two interpolants are in close agreement (better than $3$ decimal places), with the exception of the strongest-field region and spins greater than $0.9$ (regardless of eccentricity). Note that the real and imaginary components of ${\cal A}_{220}$ are interpolated separately; here we have plotted the relative difference between the complex amplitudes for visual clarity, as the components pass through zero more frequently.
@@
Recall that, in our amplitude interpolation framework (Section III.2), we opt to perform linear interpolation with respect to $a$ for the purposes of efficiency and to reduce memory usage. This is motivated by the weak relationship between mode amplitude and spin (given constant $p$ and $e$). However, our grid coordinate parametrization (which maps the curved inner surface of both the separatrix and the eccentricity taper to rectilinear coordinates) is highly non-linear near the separatrix, especially as $a\rightarrow 1$. This effect is the main contributor to the larger amplitude mismatches we observe in Fig. 9. To demonstrate this point and explore the impact of such behaviour on linear interpolation accuracy near the separatrix, we construct tricubic spline interpolants (identical in form to those applied to the forcing functions in Section III.1.1) and compare their outputs to those of the bicubic+linear scheme. For regions where the linear interpolation is a good approximation, we expect the outputs of these two interpolation methods to be in strong agreement. We consider the ${\mathcal{A}}_{220}$ mode, which is typically the most significant mode in terms of SNR contribution for most of the inspiral: we found that comparisons performed for other strongly-contributing modes yielded similar results to this case. Relative differences between the (complex) outputs of the bicubic+linear and tricubic interpolation configurations for this mode are shown in Fig. 10. We find that (except for regions of the parameter space where $|{\mathcal{A}}_{220}|\rightarrow 0$) a linear interpolation in $a$ is sufficient for mode amplitudes to be computed accurately (better than 3 digits of precision) for $p-p_{\mathrm{sep}}\gtrsim 1$. Larger relative differences occur where $|{\mathcal{A}}_{220}|\sim 0$, but as the amplitude is small here regardless of the interpolant used we do not expect this to significantly impact waveform accuracy. For lower values of $p$ and $a\gtrsim 0.9$, the relative difference grows, reaching a few tens of percent at the highest spins and lowest orbital separations. However, it is reasonable to expect that the impact of these errors on waveform accuracy will be subdominant with respect to errors of similar scale at larger orbital separations, as EMRIs typically spend only a small fraction of their number of orbital cycles in this region.
@@
For systems with particularly small mass ratios $\epsilon\lesssim\sim 10^{-6}$, which inspiral sufficiently slowly to accumulate hundreds of orbital cycles in the region $p-p_{\mathrm{sep}}<1$, the impact of these amplitude interpolation errors is likely to be more significant. As the primary focus of our waveform model is mass ratios $\epsilon\gtrsim 10^{-6}$, this is an acceptable limitation of our model (but will examine a candidate source with $\epsilon=10^{-6}$ to understand this in more detail). In Section IV.3, we will demonstrate that this reasoning holds for such systems by performing comparisons between amplitude models on the waveform level.
@@ IV.3 Waveform validation
Now that we have characterised the dominant sources of error in our trajectory and mode amplitude generation frameworks, we will now explore how these errors manifest in waveform templates produced with our model. The gold standard for understanding the impact of systematic errors is to examine them through the lens of Bayesian parameter estimation, identifying how these errors propagate to the structure of the posterior distribution (Eq. 59). However, this procedure requires millions of likelihood evaluations (and therefore hours of GPU wall-time) per comparison, and is therefore prohibitively computationally expensive for a global examination of the parameter space. Techniques for approximating linear biases in parameter estimates Miller2005 ; Flanagan:1997kp ; Cutler:2007mi (such as the approach we apply in Section V.3 for this purpose) reduce this to $\sim 10^{2}$ waveform evaluations per point in parameter space (which are required for computing waveform derivatives). However, the resulting bias vectors in 13-dimensional parameter space are challenging to interpret in the context of a random draw of sample points over the EMRI parameter space.
We instead opt to quantify the fidelity of our model in terms of the waveform mismatch $\mathcal{M}$, which is defined in terms of the noise-weighted inner product and therefore incorporates the sensitivity curve of the detector (in this case, LISA). Definitions for the inner product and mismatch are given in Appendix D. A mismatch of zero implies that two templates are identical, whereas a mismatch of one indicates that the two waveforms are orthogonal. The accuracy requirements for a waveform model to be applied in data analysis without introducing bias must necessarily be expressed relative to the SNR $\rho$ of the signal being analysed (with louder signals requiring more faithful models — see Cutler:2007mi for more discussion). A useful (albeit typically strongly conservative) criterion derived in Refs. lindblom2008model ; Chatziioannou_2017 ; purrer2020gravitational ; Toubiana:2024car relates the SNR of the template signal $s$ ($\rho_{s}$) defined in Eq. 58 to the minimum mismatch between this signal and a waveform template $h$ such that the two are indistinguishable. It states that if
@@
with $D=12$ being the number of parameters in the waveform model that are estimated during inference (cf. Table 1), then $s$ and $h$ are identical for the purposes of data analysis (at a confidence of $1\,\sigma$). Choosing what constitutes an acceptable value of $\mathcal{M}_{\mathrm{min}}$ for our model therefore requires us to specify $\rho_{s}$. For EMRIs with $\epsilon\lesssim 10^{-4}$ situated at astrophysical distances ($d_{\mathrm{L}}\gtrsim 1$ Gpc), LISA accumulates an SNR of up to $\sim 10^{3}$ over a few years of observation, given source parameters that maximise detectability (such as high MBH spin). See Table 1 for some example systems, which we will study in more detail in Section V.6. Informed by this order-of-magnitude estimate, we choose the value $\mathcal{M}_{\mathrm{min}}=10^{-5}$; this is compatible with signal SNRs of several hundreds and (considering the conservative nature of Eq. 15) should represent sufficient accuracy for the purposes of EMRI data analysis with LISA. In what follows, we stress that we do not maximise our mismatch calculations over phase, coalescence time or other extrinsic parameters (such as source orientation); as this would always improve the mismatches we would observe, our results are therefore slightly conservative. However, the impact of this choice is small for EMRIs, as these mismatches are the result of small differences between many sets of harmonic modes that (in most cases) cannot all be significantly reduced by modifying such parameters.
Figure 11: Similar to the top panel of Fig. 8, but in terms of waveform mismatches instead of orbital dephasings. We consider four waveform models built with forcing function data grids of successively coarser resolutions. All mismatches are computed with respect to waveforms generated with the finest-resolution model. The majority of mismatches decrease exponentially (in proportion to the data grid density); the red dashed histogram is obtained by extrapolating this exponential scaling to the actual resolution of our data grid. A long upper tail of mismatches is evident, which corresponds to systems with high eccentricities (see Fig. 12 and main text for discussion).
Figure 12: Mismatches between waveforms computed with models using full-resolution $(h^{\mathrm{1DS}})$ and half-resolution $(h^{\mathrm{2DS}})$ forcing function data grids (blue histogram in Fig. 11), with respect to $(e_{0},p_{0})$ (top panel) and $(e_{0},a)$ (bottom panel). The mismatches shown can be interpreted as upper bounds on the accuracy of our full-resolution model at each point in parameter space. The accuracy of our model decreases significantly for $e_{0}\gtrsim 0.85$.
@@ IV.3.1 Trajectory dephasing
Based on the results presented in Sections IV.2 and IV.1, the most significant source of systematic error in the constituent components of our waveform model (i.e., the trajectory and mode amplitude modules) is orbital dephasing due to interpolation errors in the forcing functions. We will now quantify how this dephasing propagates to the accuracy of the computed waveforms, following a similar approach as was applied in Section IV.1 but with a focus on waveform mismatch (rather than orbital dephasing). Randomly drawing $10^{4}$ points in parameter space (obtained according to the procedure detailed in Appendix C), we compute waveforms with both our (full-resolution) model and lower-resolution models with down-sampled forcing function data grids. In Fig. 11, we summarize the mismatches between the full-resolution and lower-resolution models at these points in parameter space. As in Fig. 8, a clear trend is observed, with mismatches decreasing in proportion to grid resolution. Fitting this exponential trend between grid resolution and mismatch for each data point (similarly to in Section IV.1) and extrapolating this fit to our full grid density yields the red dashed histogram in Fig. 11. Due to the poor phase accuracy of the lowest-resolution model (which regularly exhibits in excess of one radian of orbital dephasing) we do not include it when constructing this extrapolation. This should be considered a rough, order-of-magnitude estimate of the accuracy of our full-resolution model; it is not completely representative of the mismatch between our model and an error-free adiabatic model (especially for the smaller mismatches $\lesssim 10^{-8}$, as we expect other minor sources of systematic error to become important at this level). In general, we observe mismatches between waveforms from the full- and half-resolution models ($h^{\mathrm{1DS}}$ and $h^{\mathrm{2DS}}$ respectively) of less than $10^{-5}$ over the majority of the parameter space, with extrapolated mismatches improving upon this by roughly two orders of magnitude. When compared to our mismatch requirement of $10^{-5}$, these results suggest that our waveform model is highly robust over a large fraction of the EMRI parameter space.
@@
We also find a small tail of higher mismatches $\mathcal{M}\rightarrow 1$ that are easily identified when $\mathcal{M}$ is examined as a function of $(a,p,e)$, as shown in Fig. 12. Mismatch increases rapidly with respect to eccentricity beyond $e\sim 0.85$, which is not unexpected given the larger errors in the forcing functions in this region identified in Fig. 5. It is not clear whether the relatively poor performance of our model in this region is the result of data quality issues or insufficient interpolation density (or more likely, some combination of the two); regardless of the cause, these results suggest that users of our waveform model should exercise caution when choosing $e_{0}\gtrsim 0.85$. Some poorer mismatches are also observed for very high spins ($a\gtrsim 0.998$), which we attribute to the impact of insufficient grid density with respect to $a$ nearer to the separatrix (as was identified in Fig. 6). Both of these problematic regions are near the edges of our domain of validity, where the forcing functions vary rapidly, are computationally expensive and are challenging to obtain accurately. Given the resulting challenge in modeling systems with such high eccentricities and/or spins, treating these areas of the parameter space separately (e.g., by constructing separate data grids bespoke to these regions) is an idea worthy of investigation in future work.

### v2

@@ IV.2 Mode amplitude validation
We now turn to examine the faithfulness of the amplitude module of our waveform generation framework. In general, any systematic errors in the mode amplitudes $A_{\ell mn}$ should be significantly less impactful than those present in the forcing functions. Contributions from the latter grow secularly over inspiral, whereas amplitude errors remain of fixed order. For systematic biases in EMRI parameter estimation to be sufficiently small (assuming no orbital phase errors), the relative error in each mode amplitude must be smaller (in an order-of-magnitude sense) than the inverse of the SNR of the mode lindblom2008model . It is therefore important to verify that the scale of these errors is small. In this subsection, we will compare the outputs of our amplitude module with an independent dataset to quantify the impact of errors in our interpolated amplitudes. We will also explore where our choice to perform linear interpolation with respect to $a$ (as described in Section˜III.2.2) limits the accuracy our mode amplitude values when compared to spline interpolation along this dimension.
As in Section˜IV.1, we will validate our model by comparing interpolated mode amplitudes ($\mathcal{A}_{\ell mn}^{\mathrm{FEW}}$) against those produced with the GREMLIN code ($\mathcal{A}_{\ell mn}^{\mathrm{SAH}}$). As our amplitude module produces $6993$ modes at each point in parameter space, examining any particular mode in isolation will not be particularly informative with regards to the accuracy of our amplitude model as a whole (especially as the relative importance of each mode is also a function of the parameter space). Instead, in a similar vein to Ref. Isoyama:2021jjd , we will define a suitable figure of merit that encapsulates all mode amplitudes. It is similar in form to that of the waveform mismatch (which we will apply in Section˜IV.3; see Appendix˜D for details), and is written as
@@
where $\star$ denotes complex conjugation and $|\cdot|$ is the amplitude vector norm (over the indices $\ell$, $m$ and $n$). This quantity corresponds to a mismatch between two snapshot waveforms (i.e., constant in amplitude and frequency) with amplitudes $\mathcal{A}^{\mathrm{FEW}}_{\ell mn}$ and $\mathcal{A}^{\mathrm{SAH}}_{\ell mn}$, assuming that overlap terms between modes is small. While this is not generally true, the inclusion of these terms is not required for us to explore how the accuracy of our modes varies over parameter space, and their exclusion enables us to avoid performing waveform computations at this stage. This also ignores the spin-weighted spherical harmonics, as we wish to investigate differences independent of the viewing angle. The $\mathcal{A}^{\mathrm{SAH}}_{\ell mn}$ data we compare against only extends to $p\sim 10$; see Section˜E.2 for a comparison between our amplitude interpolation and PN results in the weak-field limit. In Fig.˜9, we show how $\mathcal{M}_{\mathrm{amp}}$ varies over $(p,e)$ for $a=0.8952$ (which lies between two nodes in our amplitude data grids, maximising interpolation errors). For $p\gtrsim p_{\mathrm{sep}}+1$, our interpolated amplitudes agree well with the GREMLIN data for all eccentricities, with estimated mismatches $\leq 10^{-5}$. Below this value, $\mathcal{M}_{\mathrm{amp}}$ grows rapidly with respect to $p$, reaching $\sim 10^{-3}$ at higher eccentricities for $p\gtrsim p_{\mathrm{sep}}+10^{-2}$. This suggests that our amplitude interpolation errors are significant near the separatrix. In Section˜IV.3, we investigate the impact of amplitude interpolation on the waveform level and obtain results that agree quantitatively with those presented in Fig.˜9.
@@
Figure 9: Mode amplitude (approximate) mismatch $\mathcal{M}_{\mathrm{amp}}$ (Eq.˜14) between interpolated amplitudes from our model ($\mathcal{A}_{\ell mn}^{\mathrm{FEW}}$) and those produced with the GREMLIN code ($\mathcal{A}_{\ell mn}^{\mathrm{SAH}}$). Mismatch due to amplitude interpolation error increases significantly as $p$ decreases below $\sim p_{\mathrm{sep}}+1$, and is slightly worse for higher eccentricities. This behaviour is due to our linear interpolation of amplitudes with respect to $a$ (explored in Fig.˜10).
@@
Figure 10: Relative difference between bicubic+linear (BIC) and tricubic interpolation (TRI) of the $A_{220}$ mode amplitude with respect to: $(p,e)$ with $a=0.998$ (top panel); $(p,a)$ with grid coordinate $w=0.5$ (bottom panel). The eccentricity corresponding to this value of $w$ for $a=0.998$ is shown in the top panel as a cyan dashed line; for lower spins, $e$ at $p=p_{\mathrm{sep}}+10^{-3}$ will be higher than this line (see Appendix˜B). The outputs of the two interpolants are in close agreement (better than 33 decimal places), with the exception of the strongest-field region and spins greater than $0.9$ (regardless of eccentricity). Note that the real and imaginary components of ${\cal A}_{220}$ are interpolated separately; here we have plotted the relative difference between the complex amplitudes for visual clarity, as the components pass through zero more frequently.
@@
Recall that, in our amplitude interpolation framework (Section˜III.2), we opt to perform linear interpolation with respect to $a$ for the purposes of efficiency and to reduce memory usage. This is motivated by the weak relationship between mode amplitude and spin (given constant $p$ and $e$). However, our grid coordinate parametrization (which maps the curved inner surface of both the separatrix and the eccentricity taper to rectilinear coordinates) is highly non-linear near the separatrix, especially as $a\rightarrow 1$. This effect is the main contributor to the larger amplitude mismatches we observe in Fig.˜9. To demonstrate this point and explore the impact of such behaviour on linear interpolation accuracy near the separatrix, we construct tricubic spline interpolants (identical in form to those applied to the forcing functions in Section˜III.1.1) and compare their outputs to those of the bicubic+linear scheme. For regions where the linear interpolation is a good approximation, we expect the outputs of these two interpolation methods to be in strong agreement. We consider the ${\mathcal{A}}_{220}$ mode, which is typically the most significant mode in terms of SNR contribution for most of the inspiral: we found that comparisons performed for other strongly-contributing modes yielded similar results to this case. Relative differences between the (complex) outputs of the bicubic+linear and tricubic interpolation configurations for this mode are shown in Fig.˜10. We find that (except for regions of the parameter space where $|{\mathcal{A}}_{220}|\rightarrow 0$) a linear interpolation in $a$ is sufficient for mode amplitudes to be computed accurately (better than 3 digits of precision) for $p-p_{\mathrm{sep}}\gtrsim 1$. Larger relative differences occur where $|{\mathcal{A}}_{220}|\sim 0$, but as the amplitude is small here regardless of the interpolant used we do not expect this to significantly impact waveform accuracy. For lower values of $p$ and $a\gtrsim 0.9$, the relative difference grows, reaching a few tens of percent at the highest spins and lowest orbital separations. However, it is reasonable to expect that the impact of these errors on waveform accuracy will be subdominant with respect to errors of similar scale at larger orbital separations, as EMRIs typically spend only a small fraction of their number of orbital cycles in this region.
@@
For systems with particularly small mass ratios $\epsilon\lesssim\sim 10^{-6}$, which inspiral sufficiently slowly to accumulate hundreds of orbital cycles in the region $p-p_{\mathrm{sep}}<1$, the impact of these amplitude interpolation errors is likely to be more significant. As the primary focus of our waveform model is mass ratios $\epsilon\gtrsim 10^{-6}$, this is an acceptable limitation of our model (but will examine a candidate source with $\epsilon=10^{-6}$ to understand this in more detail). In Section˜IV.3, we will demonstrate that this reasoning holds for such systems by performing comparisons between amplitude models on the waveform level.
@@ IV.3 Waveform validation
Now that we have characterised the dominant sources of error in our trajectory and mode amplitude generation frameworks, we will now explore how these errors manifest in waveform templates produced with our model. The gold standard for understanding the impact of systematic errors is to examine them through the lens of Bayesian parameter estimation, identifying how these errors propagate to the structure of the posterior distribution (Eq.˜60). However, this procedure requires millions of likelihood evaluations (and therefore hours of GPU wall-time) per comparison, and is therefore prohibitively computationally expensive for a global examination of the parameter space. Techniques for approximating linear biases in parameter estimates Miller2005 ; Flanagan:1997kp ; Cutler:2007mi (such as the approach we apply in Section˜V.3 for this purpose) reduce this to $\sim 10^{2}$ waveform evaluations per point in parameter space (which are required for computing waveform derivatives). However, the resulting bias vectors in 13-dimensional parameter space are challenging to interpret in the context of a random draw of sample points over the EMRI parameter space.
We instead opt to quantify the fidelity of our model in terms of the waveform mismatch $\mathcal{M}$, which is defined in terms of the noise-weighted inner product and therefore incorporates the sensitivity curve of the detector (in this case, LISA). Definitions for the inner product and mismatch are given in Appendix˜D. A mismatch of zero implies that two templates are identical, whereas a mismatch of one indicates that the two waveforms are orthogonal. The accuracy requirements for a waveform model to be applied in data analysis without introducing bias must necessarily be expressed relative to the SNR $\rho$ of the signal being analysed (with louder signals requiring more faithful models — see Cutler:2007mi for more discussion). A useful (albeit typically strongly conservative) criterion derived in Refs. lindblom2008model ; Chatziioannou_2017 ; purrer2020gravitational ; Toubiana:2024car relates the SNR of the template signal $s$ ($\rho_{s}$) defined in Eq.˜59 to the minimum mismatch between this signal and a waveform template $h$ such that the two are indistinguishable. It states that if
@@
with $D=12$ being the number of parameters in the waveform model that are estimated during inference (cf. Table˜1), then $s$ and $h$ are identical for the purposes of data analysis (at a confidence of $1\,\sigma$). Choosing what constitutes an acceptable value of $\mathcal{M}_{\mathrm{min}}$ for our model therefore requires us to specify $\rho_{s}$. For EMRIs with $\epsilon\lesssim 10^{-4}$ situated at astrophysical distances ($d_{\mathrm{L}}\gtrsim 1$ Gpc), LISA accumulates an SNR of up to $\sim 10^{3}$ over a few years of observation, given source parameters that maximise detectability (such as high MBH spin). See Table˜1 for some example systems, which we will study in more detail in Section˜V.6. Informed by this order-of-magnitude estimate, we choose the value $\mathcal{M}_{\mathrm{min}}=10^{-5}$; this is compatible with signal SNRs of several hundreds and (considering the conservative nature of Eq.˜15) should represent sufficient accuracy for the purposes of EMRI data analysis with LISA. In what follows, we stress that we do not maximise our mismatch calculations over phase, coalescence time or other extrinsic parameters (such as source orientation); as this would always improve the mismatches we would observe, our results are therefore slightly conservative. However, the impact of this choice is small for EMRIs, as these mismatches are the result of small differences between many sets of harmonic modes that (in most cases) cannot all be significantly reduced by modifying such parameters.
Figure 11: Similar to the top panel of Fig.˜8, but in terms of waveform mismatches instead of orbital dephasings. We consider four waveform models built with forcing function data grids of successively coarser resolutions. All mismatches are computed with respect to waveforms generated with the finest-resolution model. The majority of mismatches decrease exponentially (in proportion to the data grid density); the red dashed histogram is obtained by extrapolating this exponential scaling to the actual resolution of our data grid. A long upper tail of mismatches is evident, which corresponds to systems with high eccentricities (see Fig.˜12 and main text for discussion).
Figure 12: Mismatches between waveforms computed with models using full-resolution $(h^{\mathrm{1DS}})$ and half-resolution $(h^{\mathrm{2DS}})$ forcing function data grids (blue histogram in Fig.˜11), with respect to $(e_{0},p_{0})$ (top panel) and $(e_{0},a)$ (bottom panel). The mismatches shown can be interpreted as upper bounds on the accuracy of our full-resolution model at each point in parameter space. The accuracy of our model decreases significantly for $e_{0}\gtrsim 0.85$.
@@ IV.3.1 Trajectory dephasing
Based on the results presented in Sections˜IV.2 and IV.1, the most significant source of systematic error in the constituent components of our waveform model (i.e., the trajectory and mode amplitude modules) is orbital dephasing due to interpolation errors in the forcing functions. We will now quantify how this dephasing propagates to the accuracy of the computed waveforms, following a similar approach as was applied in Section˜IV.1 but with a focus on waveform mismatch (rather than orbital dephasing). Randomly drawing $10^{4}$ points in parameter space (obtained according to the procedure detailed in Appendix˜C), we compute waveforms with both our (full-resolution) model and lower-resolution models with down-sampled forcing function data grids. In Fig.˜11, we summarize the mismatches between the full-resolution and lower-resolution models at these points in parameter space. As in Fig.˜8, a clear trend is observed, with mismatches decreasing in proportion to grid resolution. Fitting this exponential trend between grid resolution and mismatch for each data point (similarly to in Section˜IV.1) and extrapolating this fit to our full grid density yields the red dashed histogram in Fig.˜11. Due to the poor phase accuracy of the lowest-resolution model (which regularly exhibits in excess of one radian of orbital dephasing) we do not include it when constructing this extrapolation. This should be considered a rough, order-of-magnitude estimate of the accuracy of our full-resolution model; it is not completely representative of the mismatch between our model and an error-free adiabatic model (especially for the smaller mismatches $\lesssim 10^{-8}$, as we expect other minor sources of systematic error to become important at this level). In general, we observe mismatches between waveforms from the full- and half-resolution models ($h^{\mathrm{1DS}}$ and $h^{\mathrm{2DS}}$ respectively) of less than $10^{-5}$ over the majority of the parameter space, with extrapolated mismatches improving upon this by roughly two orders of magnitude. When compared to our mismatch requirement of $10^{-5}$, these results suggest that our waveform model is highly robust over a large fraction of the EMRI parameter space.
@@
We also find a small tail of higher mismatches $\mathcal{M}\rightarrow 1$ that are easily identified when $\mathcal{M}$ is examined as a function of $(a,p,e)$, as shown in Fig.˜12. Mismatch increases rapidly with respect to eccentricity beyond $e\sim 0.85$, which is not unexpected given the larger errors in the forcing functions in this region identified in Fig.˜5. It is not clear whether the relatively poor performance of our model in this region is the result of data quality issues or insufficient interpolation density (or more likely, some combination of the two); regardless of the cause, these results suggest that users of our waveform model should exercise caution when choosing $e_{0}\gtrsim 0.85$. Some poorer mismatches are also observed for very high spins ($a\gtrsim 0.998$), which we attribute to the impact of insufficient grid density with respect to $a$ nearer to the separatrix (as was identified in Fig.˜6). Both of these problematic regions are near the edges of our domain of validity, where the forcing functions vary rapidly, are computationally expensive and are challenging to obtain accurately. Given the resulting challenge in modeling systems with such high eccentricities and/or spins, treating these areas of the parameter space separately (e.g., by constructing separate data grids bespoke to these regions) is an idea worthy of investigation in future work.